OpenStudy (anonymous):

can someone please explain 3^(2x+1)=2(5^x) ?? thank you!!!

OpenStudy (ash2326):

Is this your question? \[3^{(2x+1)}=2\times 5^x\]

OpenStudy (anonymous):

\[3^{2x+1}=2(5^x)\]

OpenStudy (anonymous):

yup!

OpenStudy (ash2326):

Do you know logarithms?

OpenStudy (ash2326):

@dydlf ?

OpenStudy (anonymous):

yup! that's our lesson too!:)

OpenStudy (ash2326):

Ok, Let's take logarithm to the base 10 both sides \[\log(3^{2x+1})=\log(2(5^x))\] Now we'll use two important properties of logarithms to simplify this \[\log (A\times B)=\log A+\log B\] \[\log (A^a)=a\times \log A\] Do you know these properties?

OpenStudy (anonymous):

yup i know those properties! :D wait i'll be back

OpenStudy (ash2326):

@dydlf could you try to simplify using these properties?

OpenStudy (anonymous):

\[\frac{ (3^{2x+1})}{ 5^x }=2\] then i multiplied the left with log base 5?!?! is that legal if ever? anyway i tried this too... i divided 2 in both sides soooo \[\frac{ 3 }{ 2 } ^{2x+1}=5^x\] theeen... i did log with base 5 soooo i ended with x at the other side then i brought the exponent 2x+1 to the left so i ended up with log 3/2 base 5=x/2x+1 ?? :( and i can't solve for x anymore :(

OpenStudy (ash2326):

Multiplying log won't help, I'll take log ( base 10) both sides \[\log(3^{2x+1})=\log(2(5^x))\] Now we'll use the properties to simplify this \[(2x+1)\times \log 3=\log 2+\log 5^{x}\]

OpenStudy (ash2326):

Do you get this? I have used these properties \[\log AB=\log A+\log B\] \[\log A^a=a\log A\]

OpenStudy (anonymous):

OOOOHHH i'm seeing something wait

OpenStudy (ash2326):

ok

OpenStudy (anonymous):

is it x= log2-log3/2log3-log5 ??

OpenStudy (ash2326):

Yes, you're right

OpenStudy (anonymous):

yay thank you!!!